Find the percent composition of ammonium hydroxide, NH4OH


1
Expert's answer
2020-12-18T08:43:46-0500

Molar mass of NH4OH = 35g/mol


atomic mass of Hydrogen = 1g ; here we have 5 hydrogen so total mass of hydrogen = 5g

% of hydrogen present = (5/35)×100 = 14.28 %


atomic mass of oxygen = 16g

% of oxygen present = (16/35)×100 = 45.72%


atomic mass of nitrogen = 14g

% of nitrogen present = (14/35) ×100 = 40%